OpenStudy (anonymous):

How do you solve 'If x/5 and x/9 are both positive integers, which of the following must also be an integer? A: x/40 B: x/25 C: x/18 D: x/15 E: x/10 i'm not sure how to solve this :c

OpenStudy (anonymous):

I would say A

OpenStudy (anonymous):

why? o:

OpenStudy (anonymous):

If x is divisible by five and by nine, then it's divisible by their product, which is 45. But 45 is 15 * 3, so if it's divisible by 45 it's certainly divisible by 15.

OpenStudy (anonymous):

ohh, so it's x/15?

OpenStudy (anonymous):

If x/5 is an integer, then we can write that x = 5*y, where y is an integer. if x/9 = 5*y/9 is an integer, than we can write y = 9z, where z is also an integer. Thus, x = 5*9*z = 45 z, where z is an integer. So x / 15 = 3 z, which must surely be an integer. Yes, to answer your question.

OpenStudy (anonymous):

thanks (: i get it